The width of a rectangle is 2 ft less than the length. The area is 15 ft^2. Find the length and the width.
The width is ___ft.
The length is ___ft.

Answer by jorel1380(3719)

l(l-2)=15
l2-2l-15=0
(l-5)(l+3)=0
The length is 5 ft.
The width is 3 ft..